# SALT: Parameter-Efficient Fine-Tuning via Singular Value Adaptation with Low-Rank Transformation

### v1

@@ 3 Experiments
Datasets. In this work, we use datasets spanning three anatomical domains: neurovascular analysis with DIAS [12], a dynamic 2D+time DSA dataset for intracranial artery segmentation in cerebrovascular disease (20 train/10 test). Retinal assessment via ROSE [14] for OCT-Angiography-based microvasculature analysis (22 train/8 test). DRIVE [22] for diabetic retinopathy screening in RGB fundus images (14 train/6 test). Cardiovascular studies using ARCADE [17] (700 train/300 test) and XRay-Angio [2] (93 train/41 test), both leveraging X-ray angiography, where the former is for region-based coronary artery disease diagnostics and the latter is for multiscale segmentation of occluded vessels. The datasets were randomly partitioned into training and testing subsets to ensure unbiased evaluation.
Implementation Details. We apply SALT to the image encoder layers of the prompt-adapted SAM with a 512 input size and text-based prompts. Following the pipeline in S-SAM [16], training uses AdamW ($\text{LR}=1\times 10^{-4}$, weight decay $=1\times 10^{-2}$), batch size 5, and 200 epochs, with a step decay scheduler for learning rate adjustment. Data augmentation includes resizing, random flips, ±$\pm 10^{\circ}$ rotations, and brightness adjustments. The total loss combines focal loss, Dice loss, and regularization loss (scaled by 0.01). Evaluation metrics include the Dice Similarity Coefficient (DSC) and 95th percentile Hausdorff Distance (HD95). Experiments were conducted on an NVIDIA RTX 4090 (24GB).
@@ 4 Results and Discussion
Model Rank %Trainable ARCADE DIAS ROSE XRay-Angio Drive Average
@@
U-Net [20] 100% / 6.5M 0.36 157.14 0.13 262.66 0.63 8.07 0.71 30.40 0.18 84.43 0.40 108.54
UNETR [5] 100% / 91M 0.48 162.44 0.31 183.97 0.63 8.96 0.74 20.62 0.46 33.58 0.52 81.91
RegUnet [20] 100% / 7.8M 0.11 247.79 0.35 172.09 0.22 58.08 0.42 79.51 0.43 72.97 0.31 126.08
SegNet [1] 100% / 29M 0.14 231.11 0.13 295.06 0.38 9.09 0.09 110.94 0.14 79.03 0.18 145.05
DeepLabV3+ [3] 100% / 45M 0.55 103.22 0.55 77.66 0.52 12.90 0.74 14.03 0.46 27.79 0.56 47.12
Segformer [24] 100% / 43M 0.53 113.98 0.57 73.64 0.56 11.21 0.75 16.55 0.48 36.98 0.58 50.47
@@
SAM w/ text prompts 0.00 - 0.00 - 0.00 - 0.04 - 0.00 - 0.01 -
S-SAM [16] 0.40% / 1M 0.76 52.44 0.67 42.36 0.65 12.41 0.75 27.56 0.70 15.81 0.71 30.12
LoRA [7] 4 0.64% / 1.5M 0.78 39.58 0.61 27.81 0.59 16.26 0.73 26.78 0.62 21.04 0.66 26.29
LoRA [7] 256 14.08% / 33.9M 0.81 36.89 0.63 34.36 0.62 14.62 0.76 24.84 0.68 19.00 0.70 25.94
SALT (ours) 4 0.46% / 1.1M 0.78 51.74 0.69 39.58 0.66 12.19 0.75 29.55 0.72 16.04 0.72 29.82
SALT (ours) 256 3.9% / 9.4M 0.81 42.50 0.71 27.35 0.67 12.64 0.77 23.67 0.75 13.20 0.74 23.87
@@
We evaluate the performance of SALT against SOTA PEFT techniques, specifically LoRA [7] and S-SAM [16]. As shown in Table 1, our experimental evaluation demonstrates the effectiveness of SAM-based PEFT methods compared to traditional deep learning approaches across five medical imaging datasets. While traditional methods like DeepLabV3+ [3] and U-Net [20] showed competitive performance on datasets such as XRay-Angio [2], they require training of all parameters. On the other hand, SALT achieves superior performance with significantly fewer trainable parameters. With rank-256 configuration, attaining an average Dice score of 0.74 while requiring only 3.9% of parameters to be trained. This represents a substantial improvement over both traditional methods and other parameter-efficient approaches like LoRA [7] and S-SAM [16]. Notably, SALT consistently achieves higher Dice scores over other methods across all datasets, achieving the highest scores on DIAS (0.71), ROSE (0.67), XRay-Angio (0.77), and Drive (0.75). The rank configuration comparison reveals that while rank-4 SALT delivers competitive results with minimal parameter overhead (0.46% trainable), the rank-256 version provides optimal performance with a modest increase in trainable parameters. Our analysis of boundary accuracy via HD95 further underscores SALT’s performance. While traditional methods exhibited high HD95 errors, such as SegNet [1] scoring 145.05 and RegUnet [20] scoring 126.08, SAM-based PEFT approaches showed improvements, with SALT achieving the lowest average HD95 ($23.87$) against LoRA[7] ($25.94$) and S-SAM [16] ($30.12$). Notably, SALT demonstrated Strong boundary adherence in challenging cases. SALT’s 13.20 HD95 on DRIVE [22] against S-SAM [16] achieving 15.81 reflects precise vessel delineation in low-contrast retinal images. This dual advantage in both Dice and HD95 confirms SALT’s ability to maintain structural accuracy and boundary precision while operating in parameter-efficient regimes.
Qualitative Analysis. As illustrated in Fig 3, the qualitative results highlight the SOTA performance of our proposed SALT method across various medical imaging datasets. Compared to traditional approaches like UNETR [5] and U-Net [20], as well as other SAM-based PEFT methods such as S-SAM [16] and LoRA [7], SALT consistently delivers more precise vessel segmentation. In the ROSE [14] dataset, SALT excels at capturing the intricate, web-like patterns of retinal vessels while maintaining structural continuity with fewer false positives. For the DIAS [12] dataset, SALT effectively segments complex branching patterns, preserving both main vessel trunks and subtle secondary branches. In X-ray angiography [2] images, SALT achieves clear vessel boundaries with fewer false positives compared to traditional methods. Additionally, in the Drive [22] dataset, it successfully segments retinal vessels despite varying widths and complex backgrounds.
@@
Figure 3: Qualitative Vessel Segmentation. The figure displays segmentation results for five datasets (Arcade [17], Rose [14], Dias [12], X-ray Angio [2], Drive [22]). Columns show the ground truth, SALT, SAM-based PEFT methods, and traditional DL models. Predicted masks are white, and false negatives are red.
Figure 4: Ablation study on SALT configurations And adaptability to SAM2. a) Evaluate model performance and efficiency across LoRA ranks, with the x-axis for parameter combinations, the y-axis for Dice scores, and circle size for trainable parameter percentage. The numbers on the x-axis represent the rank of the top singular values, which are modified using scale and shift in each mentioned layer Conv2d, MLP, and MHA. b) Compares SALT, LoRA [7], and SVD PEFTs integrated with SAM2 on two datasets of different sizes.
Extension to SAM2 [19]. We extend SALT to SAM2 and perform ablation studies comparing LoRA [7], SVD-based fine tuning, and SALT on the Hiera [21] image encoder. The SAM2 [19] image encoder utilizes a hierarchical architecture, capturing multi-scale features from input images with variable embedding dimensions ranging from low-dimensional, early-stage representations to high-dimensional, later-stage ones. This design enables LoRA to adapt effectively. In contrast, the original SAM [10] image encoder employs a backbone with constant, high-dimensional embeddings, where LoRA exhibited instability and was sometimes outperformed by S-SAM’s SVD-based fine tuning method [16]. In SAM2, however, LoRA consistently outperforms relative to SVD, while SALT achieves competitive performance to LoRA with 2.4 times fewer parameters at rank 256 and 1.8× fewer at rank 4.

### v2

@@ 3 Experiments
3.1 Datasets
In this work, we use datasets spanning three anatomical domains: neurovascular analysis with DIAS [Liu et al.(2024)Liu, Tian, Wang, Xu, Li, Li, Zhao, Tian, Pan, Deng, et al.], a dynamic 2D+time DSA dataset for intracranial artery segmentation in cerebrovascular disease (20 train/10 test). Retinal assessment via ROSE [Ma et al.(2021)Ma, Hao, Xie, Fu, Zhang, Yang, Wang, Liu, Zheng, and Zhao] for OCT-Angiography-based microvasculature analysis (22 train/8 test). DRIVE [Staal et al.(2004)Staal, Abramoff, Niemeijer, Viergever, and van Ginneken] for diabetic retinopathy screening in RGB fundus images (14 train/6 test). Cardiovascular studies using ARCADE [Popov et al.(2024)Popov, Amanturdieva, Zhaksylyk, Alkanov, Saniyazbekov, Aimyshev, Ismailov, Bulegenov, Kuzhukeyev, Kulanbayeva, Kalzhanov, Temenov, Kolesnikov, Sakhov, and Fazli] (700 train/300 test) and XRay-Angio [Cervantes-Sanchez et al.(2019)Cervantes-Sanchez, Cruz-Aceves, Hernandez-Aguirre, Hernandez-Gonzalez, and Solorio-Meza] (93 train/41 test), both leveraging X-ray angiography, where the former is for region-based coronary artery disease diagnostics and the latter is for multiscale segmentation of occluded vessels. The chosen datasets span three organ systems (brain, eye, heart), four imaging modalities (DSA, OCT-A, RGB fundus, X-ray), and both temporal and spatial resolutions. This ensures that our evaluation is robust across diverse segmentation challenges. Moreover, the diversity in dataset sizes offers a robust testbed for evaluating how different PEFT methods cope with varying dataset scales, ranging from low-resource to large-scale settings, assessing their adaptability under challenging conditions. The datasets were randomly partitioned into training and testing subsets to ensure unbiased evaluation.
3.2 Implementation Details.
We apply SALT to the image encoder layers of the prompt-adapted SAM with a 512 input size and text-based prompts. Following the pipeline in S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel], training uses AdamW ($\text{LR}=1\times 10^{-4}$, weight decay $1\times 10^{-2}$), batch size 5, and 200 epochs, with a step decay scheduler for learning rate adjustment. For simplicity and consistency across datasets, we used a single text prompt “Vein”, reflecting the primary anatomical structure of interest across all datasets and avoiding dataset-specific prompt engineering. Data augmentation includes random horizontal flips (50% probability), rotations up to $\pm 10^{\circ}$ (50% probability), saturation adjustments (20% probability), and brightness scaling with random intensity (50% probability), enhancing robustness to geometric and photometric variations. Images and masks are resized with anti-aliasing and padded to maintain consistent input size, while random cropping ensures the presence of foreground pixels during training. Input images are normalized using SAM’s expected mean and standard deviation to align with the pretrained encoder distribution. The total loss combines focal loss, Dice loss, and regularization loss (scaled by 0.01). Evaluation metrics include the Dice Similarity Coefficient (DSC) and 95th percentile Hausdorff Distance (HD95). Experiments were conducted on an NVIDIA RTX 4090 (24GB). We follow the same experimental setup to train our traditional baseline models, including a standard U-Net [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] with a 5-block encoder-decoder architecture initialized with 32 feature maps, SegNet [Badrinarayanan et al.(2017)Badrinarayanan, Kendall, and Cipolla] with its original architecture, Segformer [Xie et al.(2021)Xie, Wang, Yu, Anandkumar, Alvarez, and Luo], and the other baseline methods.
@@ 4.1 Main Results
Model Rank %Trainable ARCADE XRay-Angio DIAS ROSE DRIVE Average
@@
U-Net [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] 100% / 6.5M 0.36 157.14 0.71 30.40 0.13 262.66 0.63 8.07 0.18 84.43 0.40 108.54
UNETR [Hatamizadeh et al.(2022)Hatamizadeh, Tang, Nath, Yang, Myronenko, Landman, Roth, and Xu] 100% / 91M 0.48 162.44 0.74 20.62 0.31 183.97 0.63 8.96 0.46 33.58 0.52 81.91
RegUNet [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] 100% / 7.8M 0.11 247.79 0.42 79.51 0.35 172.09 0.22 58.08 0.43 72.97 0.31 126.08
SegNet [Badrinarayanan et al.(2017)Badrinarayanan, Kendall, and Cipolla] 100% / 29M 0.14 231.11 0.09 110.94 0.13 295.06 0.38 9.09 0.14 79.03 0.18 145.05
DeepLabV3+ [Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] 100% / 45M 0.55 103.22 0.74 14.03 0.55 77.66 0.52 12.90 0.46 27.79 0.56 47.12
Segformer [Xie et al.(2021)Xie, Wang, Yu, Anandkumar, Alvarez, and Luo] 100% / 43M 0.53 113.98 0.75 16.55 0.57 73.64 0.56 11.21 0.48 36.98 0.58 50.47
@@
SAM w/ prompts 0.00 0.04 0.00 0.00 0.00 0.01
MedSAM 0.18 115.18 0.13 80.05 0.29 168.82 0.32 131.89 0.21 153.23 0.23 129.83
S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel] 0.40% / 1M 0.76 52.44 0.75 27.56 0.67 42.36 0.65 12.41 0.70 15.81 0.71 30.12
LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] 4 0.64% / 1.5M 0.78 39.58 0.73 26.78 0.61 27.81 0.59 16.26 0.62 21.04 0.66 26.29
LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] 256 14.08% / 33.9M 0.81 36.89 0.76 24.84 0.63 34.36 0.62 14.62 0.68 19.00 0.70 25.94
SALT (ours) 4 0.46% / 1.1M 0.78 51.74 0.75 29.55 0.69 39.58 0.66 12.19 0.72 16.04 0.72 29.82
SALT (ours) 256 3.9% / 9.4M 0.81 42.50 0.77 23.67 0.71 27.35 0.67 12.64 0.75 13.20 0.74 23.87
@@
We evaluate the performance of SALT against SOTA PEFT techniques, specifically LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] and S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel]. As shown in Table 1, our experimental evaluation demonstrates the effectiveness of SAM-based PEFT methods compared to traditional deep learning approaches across five medical imaging datasets. While conventional methods like DeepLabV3+ [Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] and Vanilla U-Net, [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] showed competitive performance on datasets such as XRay-Angio [Cervantes-Sanchez et al.(2019)Cervantes-Sanchez, Cruz-Aceves, Hernandez-Aguirre, Hernandez-Gonzalez, and Solorio-Meza], they require training of all parameters. Additionally, we evaluated MedSAM[Ma et al.(2024)Ma, He, Li, Han, You, and Wang], a universal medical segmentation foundation model trained on over 1.57 million image–mask pairs, under zero-shot conditions. It achieves only an average Dice score of 0.23 with an HD95 of 129.8, demonstrating that even a highly specialized backbone benefits greatly from in-domain fine-tuning. On the other hand, SALT achieves superior performance with significantly fewer trainable parameters. With rank-256 configuration, attaining an average Dice score of 0.74 while requiring only 3.9% of parameters to be trained. This represents a substantial improvement over both traditional methods and other parameter-efficient approaches like LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] and S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel]. Notably, SALT consistently achieves higher Dice scores over other methods across all datasets, achieving the highest scores on DIAS (0.71), ROSE (0.67), XRay-Angio (0.77), and DRIVE (0.75). The rank configuration comparison reveals that while rank-4 SALT delivers competitive results with minimal parameter overhead (0.46% trainable), the rank-256 version provides optimal performance with a modest increase in trainable parameters.
Table 1 quantifies the parameter and computational efficiency by comparing the number of trainable parameters for each method, providing a direct measure of tuning overhead and runtime cost. PEFT methods achieve state-of-the-art performance while significantly reducing the number of trainable parameters and computational demands. For instance, LoRA at rank 256 tunes 33.9M parameters (14%), whereas SALT at the same rank requires only 9.4M (3.9%). Both SALT and LoRA reduce overhead further to approximately 0.5% at rank 4. Furthermore, LoRA’s GFLOPsGFLOPs. We measure per-sample MACs (eval, no-grad) and report GFLOPs using $1\,\mathrm{MAC}=2\,\mathrm{FLOPs}$ ($\mathrm{GFLOPs}=2\,\mathrm{MACs}/10^{9}$). Adapter overheads are computed by our custom hooks. grow roughly linearly with rank. At rank 4, it adds 1.24 GFLOPs while at rank 256 it adds 79.2 GFLOPs, a x64 jump matching the rank increase. By contrast, SALT is nearly rank-independent because the weight recomposition step dominates. With (0.46%) trainable parameters, SALT adds 22.88 GFLOPs. At a higher rank of 256 and (3.9%) trainable parameters, this increases only by 0.64 GFLOPs to reach 23.52 GFLOPs. We exclude the one-off SVD factorization from per-sample GFLOPs, since it is incurred only once. For completeness, the total one-time setup across all SVD/SALT layers is 871.8 GFLOPs, which amortizes to $\approx 4.36$ GFLOPs per epoch over 200 epochs.
Performance gains also correlate with dataset size. LoRA tends to perform well on larger datasets like ARCADE [Popov et al.(2024)Popov, Amanturdieva, Zhaksylyk, Alkanov, Saniyazbekov, Aimyshev, Ismailov, Bulegenov, Kuzhukeyev, Kulanbayeva, Kalzhanov, Temenov, Kolesnikov, Sakhov, and Fazli] and XRay-Angio[Cervantes-Sanchez et al.(2019)Cervantes-Sanchez, Cruz-Aceves, Hernandez-Aguirre, Hernandez-Gonzalez, and Solorio-Meza] but experiences degradation on smaller datasets. In contrast, SALT consistently delivers strong and stable performance across all dataset sizes, demonstrating high robustness, especially in low-resource scenarios. This can be attributed to SALT’s targeted adaptation, where scale and shift-guided updates to the top singular values preserve useful pre-trained features, serving as a built-in regularizer that prevents overfitting on small datasets while residual low-rank updates simultaneously enhance model capacity.
Analysis of boundary accuracy via HD95 further underscores SALT’s performance. While traditional methods exhibited high HD95 errors, such as SegNet [Badrinarayanan et al.(2017)Badrinarayanan, Kendall, and Cipolla] scoring 145.05 and RegUnet [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] scoring 126.08, SAM-based PEFT approaches showed improvements, with SALT achieving the lowest average HD95 ($23.87$) against LoRA[Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] ($25.94$) and S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel] ($30.12$). Notably, SALT demonstrated strong boundary adherence in challenging cases. SALT’s 13.20 HD95 on DRIVE [Staal et al.(2004)Staal, Abramoff, Niemeijer, Viergever, and van Ginneken] against S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel] achieving 15.81 reflects precise vessel delineation in low-contrast retinal images. This dual advantage in both Dice and HD95 confirms SALT’s ability to maintain structural accuracy and boundary precision while operating in parameter-efficient regimes.
Explaining the gain in performance, and drawing from the Eckart–Young–Mirsky theorem [Eckart and Young(1936)], SALT can be interpreted as adjusting the most informative components of the weight matrices while bounding operator-norm deviations. This would help maintain stability during adaptation. From a domain adaptation perspective, these updates would reduce the covariate shift between source (natural image) and target (medical) distributions by aligning feature covariances in the dominant subspace, while the residual low-rank updates would contribute to conditional alignment.
A closer inspection reveals that boundary improvements do not always scale proportionally with Dice despite consistent gains (Table 1). We observe cases where boundary precision (HD95) does not improve at the same rate. For example, on ARCADE, SALT with rank 256 matches LoRA’s Dice score (0.81) but has a higher HD95 (42.50 vs. 36.89). At the lower rank of 4, SALT improves Dice on DIAS (0.69 vs. 0.61) yet underperforms on HD95 (39.58 vs. 27.81). These observations might suggest that SALT’s dominant-singular-value rescaling and low-rank residuals may prioritize region overlap while occasionally under-capturing high-frequency boundary cues, particularly at lower ranks.
4.2 Qualitative Analysis
@@
Figure 3: Qualitative Analysis. The figure displays segmentation results for five datasets (Arcade [Popov et al.(2024)Popov, Amanturdieva, Zhaksylyk, Alkanov, Saniyazbekov, Aimyshev, Ismailov, Bulegenov, Kuzhukeyev, Kulanbayeva, Kalzhanov, Temenov, Kolesnikov, Sakhov, and Fazli], ROSE [Ma et al.(2021)Ma, Hao, Xie, Fu, Zhang, Yang, Wang, Liu, Zheng, and Zhao], DIAS [Liu et al.(2024)Liu, Tian, Wang, Xu, Li, Li, Zhao, Tian, Pan, Deng, et al.], XRay-Angio [Cervantes-Sanchez et al.(2019)Cervantes-Sanchez, Cruz-Aceves, Hernandez-Aguirre, Hernandez-Gonzalez, and Solorio-Meza], DRIVE [Staal et al.(2004)Staal, Abramoff, Niemeijer, Viergever, and van Ginneken]). Columns show the ground truth, SALT, SAM-based PEFT methods, and traditional DL models. Predicted masks are white, and false negatives are red.
As illustrated in Fig. 3, the qualitative results highlight the SOTA performance of our proposed SALT method across various medical imaging datasets. Compared to traditional approaches like UNETR [Hatamizadeh et al.(2022)Hatamizadeh, Tang, Nath, Yang, Myronenko, Landman, Roth, and Xu] and U-Net [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox], as well as other SAM-based PEFT methods such as S-SAM [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel] and LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen], SALT consistently delivers more precise vessel segmentation. In the ROSE [Ma et al.(2021)Ma, Hao, Xie, Fu, Zhang, Yang, Wang, Liu, Zheng, and Zhao] dataset, SALT excels at capturing the intricate, web-like patterns of retinal vessels while maintaining structural continuity with fewer false positives. For the DIAS [Liu et al.(2024)Liu, Tian, Wang, Xu, Li, Li, Zhao, Tian, Pan, Deng, et al.] dataset, SALT effectively segments complex branching patterns, preserving both main vessel trunks and subtle secondary branches. In X-ray angiography [Cervantes-Sanchez et al.(2019)Cervantes-Sanchez, Cruz-Aceves, Hernandez-Aguirre, Hernandez-Gonzalez, and Solorio-Meza] images, SALT achieves clear vessel boundaries with fewer false positives compared to traditional methods. Additionally, in the DRIVE [Staal et al.(2004)Staal, Abramoff, Niemeijer, Viergever, and van Ginneken] dataset, it successfully segments retinal vessels despite varying widths and complex backgrounds.
4.3 SALT with SAM2
Figure 4: (a) Model performance and parameter efficiency across different LoRA and SALT ranks: x-axis shows parameter combinations for different layers, y-axis shows the Dice score, and the circle size indicates the percentage of trainable parameters. (b) Comparison of SALT, LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen], and SVD-based PEFT on SAM2 over two datasets of varying sizes.
We further evaluate SALT on SAM2, comparing with LoRA [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen], and SVD-based fine tuning on the Hiera [Ryali et al.(2023)Ryali, Hu, Bolya, Wei, Fan, Huang, Aggarwal, Chowdhury, Poursaeed, Hoffman, Malik, Li, and Feichtenhofer] image encoder. The SAM2 [Ravi et al.(2024)Ravi, Gabeur, Hu, et al.] image encoder utilizes a hierarchical architecture, capturing multi-scale features from input images with variable embedding dimensions ranging from low-dimensional, early-stage representations to high-dimensional, later-stage ones. This design enables LoRA to adapt effectively. In contrast, the original SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] image encoder employs a backbone with constant, high-dimensional embeddings, where LoRA exhibited instability and was sometimes outperformed by S-SAM’s SVD-based fine-tuning method [Paranjape et al.(2024)Paranjape, Sikder, Vedula, and Patel]. In SAM2, however, LoRA consistently outperforms relative to SVD, while SALT achieves competitive performance to LoRA with 2.4 times fewer parameters at rank 256 and 1.8× fewer at rank 4. According to the table shown in figure 4, SALT, at a LoRA rank configuration of 4, achieves a Dice score of 0.804 on the ARCADE [Popov et al.(2024)Popov, Amanturdieva, Zhaksylyk, Alkanov, Saniyazbekov, Aimyshev, Ismailov, Bulegenov, Kuzhukeyev, Kulanbayeva, Kalzhanov, Temenov, Kolesnikov, Sakhov, and Fazli] dataset with only 0.5 million trainable parameters, substantially reducing the parameter footprint compared to LoRA’s 0.9 million parameters. When scaled to a higher rank of 256, SALT attains Dice scores of 0.822 and 0.677 on the ARCADE [Popov et al.(2024)Popov, Amanturdieva, Zhaksylyk, Alkanov, Saniyazbekov, Aimyshev, Ismailov, Bulegenov, Kuzhukeyev, Kulanbayeva, Kalzhanov, Temenov, Kolesnikov, Sakhov, and Fazli] and ROSE [Ma et al.(2021)Ma, Hao, Xie, Fu, Zhang, Yang, Wang, Liu, Zheng, and Zhao] datasets respectively, closely matching the performance of LoRA (0.824 and 0.679) despite requiring less than half the number of trainable parameters (19.7 million versus 44.6 million). for boundary analysis, SALT achieves superior boundary segmentation on the ROSE dataset as indicated by the lowest HD95 value of 11.63, outperforming both LoRA and the baseline SVD fine-tuning. Collectively, these results show that SALT offers an effective balance between parameter efficiency and segmentation accuracy for the SAM2 architecture, enabling precise and resource-efficient fine-tuning through optimized parameter allocation.
4.4 Ablation Study